OpenStudy (anonymous):

Find the area of the region bounded by the graphs of y= \sqrt{x} and y= -x - 1 between x=1 and x=4.

OpenStudy (campbell_st):

so graph the functions use https://www.desmos.com/calculator but you are basically find the area between  the area below the x axis needs absolute values... so its basically the sum of 2 areas 1 above and 1 below the x asxis so its \[A = \int\limits_{1}^{4} (x^{\frac{1}{2}}) dx + \left| \int\limits_{1}^{4} (-x -1)dx \right|\] hope it helps